The coefficient matrix of any linear system must be square matrix.
True or false with correct explanation
1
Expert's answer
2020-10-13T14:27:52-0400

It's false.


In general, a system with m linear and n unknowns can be written as

"a_{11}x_1 + a_{12}x_2 + ... + a_{1n}x_{n} = b_1"

"a_{21}x_1 + a_{22}x_2 + ... + a_{2n}x_{n} = b_2"

"\vdots"

"a_{m1}x_1 + a_{m2}x_2 + ... + a_{mn}x_{n} = b_m"


where "x_1, x_2 ...x_n" are the unknowns and the numbers "a_{11}, a_{12}, ... a_{mn}" are the coefficients of the system. So, the coefficient matrix is the mxn matrix with the coefficient "\displaystyle a_{ij}" as the (i,j)-th entry:


"\begin{pmatrix}
a_{11} & a_{12} \space \space \space \dots & a_{1n}\\
a_{21} & a_{22} \space \space \space \dots & a_{2n}\\
&\vdots\\
a_{m1} & a_{m2} \space \space \space \dots & a_{mn}\\
\end{pmatrix}"